# AHCPTQ: Accurate and Hardware-Compatible Post-Training Quantization for Segment Anything Model

### v1

@@ Abstract
The Segment Anything Model (SAM) has demonstrated strong versatility across various visual tasks. However, its large storage requirements and high computational cost pose challenges for practical deployment. Post-training quantization (PTQ) has emerged as an effective strategy for efficient deployment, but we identify two key challenges in SAM that hinder the effectiveness of existing PTQ methods: the heavy-tailed and skewed distribution of post-GELU activations, and significant inter-channel variation in linear projection activations. To address these challenges, we propose AHCPTQ, an accurate and hardware-efficient PTQ method for SAM. AHCPTQ introduces hardware-compatible Hybrid Log-Uniform Quantization (HLUQ) to manage post-GELU activations, employing log2 quantization for dense small values and uniform quantization for sparse large values to enhance quantization resolution. Additionally, AHCPTQ incorporates Channel-Aware Grouping (CAG) to mitigate inter-channel variation by progressively clustering activation channels with similar distributions, enabling them to share quantization parameters and improving hardware efficiency. The combination of HLUQ and CAG not only enhances quantization effectiveness but also ensures compatibility with efficient hardware execution. For instance, under the W4A4 configuration on the SAM-L model, AHCPTQ achieves 36.6% mAP on instance segmentation with the DINO detector, while achieving a 7.89×7.89$\times$7.89 × speedup and 8.64×8.64$\times$8.64 × energy efficiency over its floating-point counterpart in FPGA implementation. Code will be available upon acceptance.
@@ 3.2 Challenges of SAM Quantization
The first challenge arises from the heavy-tailed and skewed distribution of post-GELU activations. As depicted in Fig. 1(a), over 90% of activations are densely concentrated within -0.2 to 0, whereas a sparser but evenly distributed subset of large values, critical for inference accuracy, extends from 0 to 0.8. As illustrated in Fig. 2 (left panel), existing log2 and uniform quantizers fail to effectively handle this distribution. The log2 quantizer efficiently captures small, densely clustered values by allocating most of its quantization grids to this range. However, its exponentially spaced grid structure results in insufficient representation density for larger values, leading to high quantization errors in the upper range. On the other hand, the uniform quantizer provides consistent resolution across the full range, making it better suited for large, evenly spaced values, but it lacks sufficient resolution for small activations, introducing substantial quantization errors. This problem is further exacerbated as bit-width decreases, particularly in ultra-low-bit settings, where the limited number of grids imposes severe constraints. While non-uniform quantizers could offer a more suitable alternative, their hardware complexity makes deployment challenging [35, 3, 9]. To effectively address this issue, an efficient, hardware-compatible quantizer that captures both small and large activations is required.
@@
The second challenge arises from high inter-channel variation, particularly in the activations of Query/Key/Value linear projections in the attention module and Linear projections in the MLP module [^1] A comprehensive analysis is provided in Appendix A.. These variations originate from LayerNorm operations and the unique activation distributions of the SAM mask decoder. As shown in Fig. 1(b), activation ranges in the Token-to-Image Value linear projection exhibit significant inter-channel disparities, making per-tensor quantization ineffective due to the challenge of finding a single optimal scale factor and zero point [1]. A large scale factor, selected to accommodate high-range channels, leads to reduced quantization granularity for low-range channels, often causing them to be rounded to zero. Conversely, a small scale factor, optimized for low-range channels, results in severe clipping in high-range channels, leading to significant information loss [43]. Moreover, while a zero point could compensate for activation skewness, the varying inter-channel median values prevent a single zero point from being optimal for all channels. While per-channel quantization can effectively mitigate quantization errors, it introduces considerable hardware inefficiencies. As depicted in Fig. 4, transferring per-channel quantization parameters between DRAM and compute units incurs a memory access overhead of up to several kB per layer [10, 36]. Storing these parameters on-chip can eliminate the transfer cost, but at the expense of a significantly larger memory footprint, requiring tens of thousands of on-chip registers, thereby increasing chip area utilization [^2] A detailed benchmarking is provided in Appendix B.. To address this trade-off, a granularity-aware quantization approach with hardware co-optimization is essential for balancing quantization accuracy and deployment efficiency.
@@ 3.3.2 Channel-Aware Grouping
To balance quantization granularity with hardware efficiency, we propose CAG, which achieves accuracy comparable to per-channel quantization while significantly reducing parameter overhead, thereby enabling efficient on-chip deployment. The fundamental concept of CAG is to progressively group channels with similar activation distributions and assign them shared quantization parameters. As outlined in Algorithm 1, the process begins with quantization parameter initialization (scales and zero points) for each channel via model calibration, treating each channel as an independent group initially. Next, block-wise reconstruction is applied to refine both quantization parameters and weights, minimizing quantization error as formulated in Eq. 5, ensuring alignment with each group’s distribution characteristics. At designated milestones, channels with similar quantization parameters are progressively clustered, and the resulting centroids are adopted as shared quantization parameters for all channels within a group. This iterative process continues until the target group count is reached. The right panel of Fig. 2 illustrates an example of this grouping process for linear projection activations in Token-to-Image Value cross-attention, demonstrating how CAG effectively reduces the number of groups while preserving quantization performance. With a group number of 4, the edge accelerator can minimize quantization parameter overhead, reducing either data transmission or on-chip storage by 99.7%, as shown in Fig. 4. In AHCPTQ, we adopt on-chip storage for quantization parameters, requiring only 144 registers to store the scales and zero points of these 4 groups under 4-bit quantization. As depicted in Fig. 6, CAG maintains accuracy comparable to per-channel quantization, offering high hardware efficiency while significantly enhancing model performance.
@@
3.4 Hardware Co-optimization
In this subsection, we briefly describe the strategy of hardware co-optimization for the proposed HLUQ and CAG. We provide a detailed implementation and illustration of the co-designed FPGA hardware in Appendix C.
HLUQ. To implement HLUQ, a dual processing-element (PE) architecture is employed to facilitate matrix multiplication within the HLUQ quantizer. The uniform quantization branch utilizes a standard integer multiplier and an adder tree, while the log2 quantization branch employs a bit-shifter and a decimal adder tree. The grid ratio between the two quantizers follows the $2^{-n}$ rule, ensuring that the MSBs of quantized values function as label bits. These label bits enable the accelerator to efficiently route quantized values to their respective PEs. Finally, the outputs from both branches are merged during dequantization, producing the final result.
CAG. To efficiently implement CAG, a small set of on-chip registers is used to store quantization parameters, and a customized quantizer logic is designed. Specifically, weights and activations are reordered by channel indices, ensuring that grouped channels are sequentially aligned. Weights are pre-processed offline, while activations are reordered on-chip. This reordering is integrated into the quantization and dequantization processes within linear projection computations. By employing counter-based logic for parameter switching, the design effectively minimizes hardware complexity while maintaining computational efficiency.
@@ 4.3 Ablation Studies
Dependence on Group Number. To analyze the effect of group number on SAM’s performance in CAG, we evaluate mAP as the group number varies from 2 to 32 in YOLOX and H-Deformable-DETR under the W4A4 configuration. The results, presented in Fig. 6, also include per-tensor and per-channel configurations for reference. For SAM-L and SAM-H models, mAP increases sharply and saturates at a group number of 2, with slight fluctuations and marginal improvements closer to the per-channel configuration. For the SAM-B model, the turning point starts at a group number of 4, with a larger performance gap of approximately 3% compared to the per-channel configuration. Considering the need for a consistent configuration across models while minimizing hardware overhead, we choose a group number of 4 in the AHCPT framework.

### v3

@@ Abstract
The Segment Anything Model (SAM) has demonstrated strong versatility across various visual tasks. However, its large storage requirements and high computational cost pose challenges for practical deployment. Post-training quantization (PTQ) has emerged as an effective strategy for efficient deployment, but we identify two key challenges in SAM that hinder the effectiveness of existing PTQ methods: the heavy-tailed and skewed distribution of post-GELU activations, and significant inter-channel variation in linear projection activations. To address these challenges, we propose AHCPTQ, an accurate and hardware-efficient PTQ method for SAM. AHCPTQ introduces hardware-compatible Hybrid Log-Uniform Quantization (HLUQ) to manage post-GELU activations, employing log2 quantization for dense small values and uniform quantization for sparse large values to enhance quantization resolution. Additionally, AHCPTQ incorporates Channel-Aware Grouping (CAG) to mitigate inter-channel variation by progressively clustering activation channels with similar distributions, enabling them to share quantization parameters and improving hardware efficiency. The combination of HLUQ and CAG not only enhances quantization effectiveness but also ensures compatibility with efficient hardware execution. For instance, under the W4A4 configuration on the SAM-L model, AHCPTQ achieves 36.6% mAP on instance segmentation with the DINO detector, while achieving a 7.89×7.89$\times$7.89 × speedup and 8.64×8.64$\times$8.64 × energy efficiency over its floating-point counterpart in FPGA implementation. Code is available at https://github.com/Keio-CSG/AHCPTQ
@@ 3.2 Challenges of SAM Quantization
The first challenge arises from the heavy-tailed and skewed distribution of post-GELU activations. As depicted in Fig. 1(a), over 90% of activations are densely concentrated within -0.2 to 0, whereas a sparser but evenly distributed subset of large values, critical for inference accuracy, extends from 0 to 0.8. As illustrated in Fig. 2 (left panel), existing log2 and uniform quantizers fail to effectively handle this distribution. The log2 quantizer efficiently captures small, densely clustered values by allocating most of its quantization grids to this range. However, its exponentially spaced grid structure results in insufficient representation density for larger values, leading to high quantization errors in the upper range. On the other hand, the uniform quantizer provides consistent resolution across the full range, making it better suited for large, evenly spaced values, but it lacks sufficient resolution for small activations, introducing substantial quantization errors. This problem is further exacerbated as bit-width decreases, particularly in ultra-low-bit settings, where the limited number of grids imposes severe constraints. Although non-uniform quantizers may provide a better alternative, their hardware complexity poses deployment challenges [36, 3, 9] and often necessitates substantial QAT-based retraining [15, 40]. To effectively address this issue, an efficient, hardware-compatible quantizer that captures both small and large activations tailored for PTQ is required.
@@
The second challenge arises from high inter-channel variation, particularly in the activations of Query/Key/Value linear projections in the attention module and Linear projections in the MLP module. These variations originate from LayerNorm operations and the unique activation distributions of the SAM mask decoder. As shown in Fig. 1(b), activation ranges in the Token-to-Image Value linear projection exhibit significant inter-channel disparities, making per-tensor quantization ineffective due to the challenge of finding a single optimal scale factor and zero point [1]. A large scale factor, selected to accommodate high-range channels, leads to reduced quantization granularity for low-range channels, often causing them to be rounded to zero. Conversely, a small scale factor, optimized for low-range channels, results in severe clipping in high-range channels, leading to significant information loss [46]. Moreover, while a zero point could compensate for activation skewness, the varying inter-channel median values prevent a single zero point from being optimal for all channels. While per-channel quantization can effectively mitigate quantization errors, it introduces considerable hardware inefficiencies. As depicted in Fig. 4, transferring per-channel quantization parameters between DRAM and compute units incurs a memory access overhead of up to several kB per layer [10, 37]. Storing these parameters on-chip can eliminate the transfer cost, but at the expense of a significantly larger memory footprint, requiring tens of thousands of on-chip registers, thereby increasing chip area utilization. To address this trade-off, a granularity-aware quantization approach with hardware co-optimization is essential for balancing quantization accuracy and deployment efficiency.
@@ 3.3.2 Channel-Aware Grouping
To balance quantization granularity with hardware efficiency, we propose CAG, which achieves accuracy comparable to per-channel quantization while significantly reducing parameter overhead, thereby enabling efficient on-chip deployment. Unlike static per-group quantization [42], the fundamental concept of CAG is to progressively group channels with similar activation distributions and assign them shared quantization parameters. As outlined in Algorithm 1, the process begins with quantization parameter initialization (scales and zero points) for each channel via model calibration, treating each channel as an independent group initially. Next, block-wise reconstruction is applied to refine both quantization parameters and weights, minimizing quantization error as formulated in Eq. 5, ensuring alignment with each group’s distribution characteristics. At designated milestones, channels with similar quantization parameters are progressively clustered, and the resulting centroids are adopted as shared quantization parameters for all channels within a group. This iterative process continues until the target group count is reached. The right panel of Fig. 2 illustrates an example of this grouping process for linear projection activations in Token-to-Image Value cross-attention, demonstrating how CAG effectively reduces the number of groups while preserving quantization performance. With a group number of 4, the edge accelerator can minimize quantization parameter overhead, reducing either data transmission or on-chip storage by 99.7%, as shown in Fig. 4. In AHCPTQ, we adopt on-chip storage for quantization parameters, requiring only 144 registers to store the scales and zero points of these 4 groups under 4-bit quantization. As depicted in Fig. 6, CAG maintains accuracy comparable to per-channel quantization, offering high hardware efficiency while significantly enhancing model performance.
3.4 Hardware Co-optimization
In this subsection, we briefly describe the strategy of hardware co-optimization for the proposed HLUQ and CAG. To implement HLUQ, a dual processing-element (PE) architecture is employed to facilitate matrix multiplication within the HLUQ quantizer. The uniform quantization branch utilizes a standard integer multiplier and an adder tree, while the log2 quantization branch employs a bit-shifter and a decimal adder tree. The grid ratio between the two quantizers follows the $2^{-n}$ rule, ensuring that the MSBs of quantized values function as label bits. These label bits enable the accelerator to efficiently route quantized values to their respective PEs. Finally, the outputs from both branches are merged during dequantization, producing the final result. To efficiently implement CAG, a small set of on-chip registers is used to store quantization parameters, and a customized quantizer logic is designed. Specifically, weights and activations are reordered by channel indices, ensuring that grouped channels are sequentially aligned. Weights are pre-processed offline, while activations are reordered on-chip. This reordering is integrated into the quantization and dequantization processes within linear projection computations. By employing counter-based logic for parameter switching, the design effectively minimizes hardware complexity while maintaining computational efficiency.
@@ 4.3 Ablation Studies
Dependence on Group Number. To analyze the effect of group number on SAM’s performance in CAG, we evaluate mAP as the group number varies from 2 to 32 in YOLOX and H-Deformable-DETR under the W4A4 configuration. The results, presented in Fig. 6, also include per-tensor and per-channel configurations for reference. For SAM-L and SAM-H models, mAP increases sharply and saturates at a group number of 2, with slight fluctuations and marginal improvements closer to the per-channel configuration. For the SAM-B model, the turning point starts at a group number of 4, with a larger performance gap of approximately 3% compared to the per-channel configuration. Considering the need for a consistent configuration across models while minimizing hardware overhead, we choose a group number of 4 in the AHCPTQ framework.
@@ 6 Conclusion
Acknowledgments
This research was supported in part by the JST CREST JPMJCR21D2 including AIP challenge program, Japan, JSPS Kakenhi 23H00467, Futaba Foundation, Asahi Glass Foundation, Telecommunications Advancement Foundation, and JST Doctoral Program Student Support Project. We also thank the anonymous reviewers and area chairs for their constructive feedback.
@@ Appendix D Experiment on Vision Transformers
Appendix E Parameter Sensitivity Analysis
To initialize $\hat{b}$, $s_{1}$, and $s_{2}$ for subsequent optimization in HLUQ, we perform a grid search over two parameters, $\alpha$ and $\beta$. We extend the search to $\alpha\in\{0.1,0.2,\ldots,0.9\}$ and $\beta\in\{\frac{1}{8},\frac{1}{4},\ldots,\frac{7}{8}\}$ on SAM-B with YOLOX. The resulting W4A4/W5A5/W6A6 mAP scores of 13.1/32.0/35.3 align with the reported values as shown in Table 5, indicating that quantization performance is empirically insensitive to these parameters and a limited subset suffices.
Table 5: Parameter sensitivity analysis on SAM-B with YOLOX.
Method W4A4 W5A6 W6A6
Default 13.4 31.8 35.4
Extended Search 13.1 32.0 35.3
Appendix F Generalizing CAG to Weight Grouping
To ensure a fair comparison with existing baselines, we apply per-channel quantization to model weights. However, CAG can also be extended to weight quantization, potentially broadening its applicability. To evaluate this, we apply CAG with 32 groups to all model weights on SAM-B with YOLOX. The resulting quantization performance is comparable to the per-channel baseline as shown in Table 6, demonstrating CAG’s generalizability to model weights.
Table 6: Performance analysis of applying CAG with 32 groups to the model weights of SAM-B using YOLOX.
Method W4A4 W5A6 W6A6
Default 13.4 31.8 35.4
+ Weight Grouping 12.3 30.8 34.3
Appendix G Comparison of Visualization Results
Figure 15: Qualitative comparison of segmentation masks generated by different quantization methods on SAM-B with YOLOX. Our AHCPTQ closely matches the floating-point reference, significantly outperforming other baselines.
We further provide visualization results on W4A4 quantization of SAM-H using YOLOX, comparing AHCPTQ with existing quantization methods, including BRECQ [17], QDrop [39], and PTQ4SAM [29], as illustrated in Fig. 15. Qualitatively, our AHCPTQ consistently generates segmentation masks closely resembling those obtained from the original floating-point model, preserving finer structural details and accurate object boundaries. In contrast, the masks produced by other PTQ methods exhibit notable degradations. For instance, QDrop and PTQ4SAM often fail to capture intricate object contours, resulting in coarse segmentation masks and inaccurate boundary delineations. BRECQ, while generally performing better than QDrop and PTQ4SAM, still experiences noticeable detail loss and fragmented segmentation in complex regions. Overall, the visualization outcomes clearly demonstrate that AHCPTQ effectively addresses the quantization challenges inherent in SAM, providing superior segmentation quality comparable to the floating-point baseline, thus confirming its effectiveness and robustness in practical low-bit deployment scenarios.